OpenStudy (anonymous):

How many permutations are there for the letters in the word LOLLIPOP? 1680 6720 10080 40320 i think i am missing a step or not because i think is d can you check my answer

OpenStudy (studygurl14):

Hold on while i check...

OpenStudy (anonymous):

OK it is my last question and last test after this i can go on spring break

OpenStudy (perl):

you have repeated letters, so you have to use a modified rule of n!

OpenStudy (perl):

divide by the number of repeats , factorial

OpenStudy (studygurl14):

yeah I was wondering about that @perl . I didn't know if they were counting repeats or not

OpenStudy (anonymous):

i know i just want to double check my answer

OpenStudy (studygurl14):

D is incorrect @onequestionatatime

OpenStudy (perl):

$$ \Large \frac{n!} {n_1! ~ n_2! ~... n_k! } ~ , where~ n_1 + n_2 + ... n_k = n $$

OpenStudy (studygurl14):

Like @perl said, you need to take that number you got, and divide it by the number of repeats. Notice in LOLLIPOP L repeats three times, O twice, and P twice

OpenStudy (studygurl14):

Does that make sense @onequestionatatime ?

OpenStudy (anonymous):

yes

OpenStudy (studygurl14):

Okay. :)

OpenStudy (studygurl14):

Make sure that the repeats have the ! thingie applied to it (forgot what it's called)

OpenStudy (perl):

$$ \Large \frac{8!} {3! ~ 2!~ 2!~1! } $$

OpenStudy (perl):

factorial :)

OpenStudy (anonymous):

yea i am doing the math right now

OpenStudy (anonymous):

i think i am doing this wrong or i got i right 3665.45

OpenStudy (anonymous):

i had a feeling

OpenStudy (perl):

can you get the denominator, that part is easy

OpenStudy (perl):

easier*

OpenStudy (studygurl14):

You already have the numerator, cause that's answer choice D

OpenStudy (anonymous):

11

OpenStudy (studygurl14):

No. Remember what 3! = 3 x 2 x 1

OpenStudy (anonymous):

is the answer 6720

OpenStudy (anonymous):

you divide D bye 6

OpenStudy (perl):

$$ \Large \frac{8!} {3! ~ 2!~ 2!~1! }=\Large \frac{8\times 7 \times6 \times5 \times4 \times3 \times2 \times1} {(3 \times 2 \times1)\cdot (2\times1)\cdot(2\times1) \cdot1 } $$

OpenStudy (perl):

the denominator is 6 x 2 x 2

OpenStudy (anonymous):

answer i got it this time it is A

OpenStudy (anonymous):

1680